Two angles are complimentary The sum of the measure of the first angle and one-fourth the second angle is 88.50. Find the measures of the angles.
Answer by mananth(16949)

let one angle be x
the other will be 90-x
..
x+1/4(90-x)=88.5
4x+90-x / 4 =88.5
3x+90=4*88.5
3x+90=354
3x=354-90
3x=264
x=264/3
x=88 deg
the other angle = 2 deg